A mineral of the type FeSO4.xH2O contains 37.2% of water.Complete the calculation to determine x.mass of one mole of H2O = 18gmass of water in 100g of FeSO4.xH2O = 37.2gnumber of moles of H2O in 100g of FeSO4.xH2O = .................mass of FeSO4 in 100g of FeSO4.xH2O = .................g
1
Expert's answer
2020-04-15T01:51:01-0400

number of moles of H2O in 100g of FeSO4.xH2O = m/M=37.2/18=2.066 mole

mass of FeSO4 in 100g of FeSO4.xH2O = 100-37.2=62.8g

number of moles of FeSO4 in 100g of FeSO4.xH2O=62.8/152=0.41 mole

n(FeSO4):n(H2O)=0.41:2.066=1:5 =>x=5

FeSO4•5H2O